# Best practices in Quantum Monte Carlo for metal catalysis: CO hydrolysis on Pt(111)

### v2

@@ Abstract
Many scientists are now researching hydrogen production since it affords a clean, sustainable alternative to fossil fuel [4]. Research work and prototyping various aspects of hydrogen power continues to advance.
Over 90 % of all chemical manufacture uses a solid catalyst. This work describes catalytic selective hydrogen production (via water-gas shift, hydration of pre-adsorbed CO) optimising reactant structure on a model catalyst. Focus is on partial O-H bond dissociation after carbon-monoxide is co-adsorbed with water at Pt(111). Finally, hydrogen gas is given off. Many chemical reactions involve bond-dissociation. This process is often the key to rate-limiting reaction steps at solid surfaces. Bond-breaking is poorly described by Hartree-Fock and DFT methods, our embedded active site approach is used. We showcase Quantum Monte Carlo (QMC) methodology using the ground-state Slater Determinant of a simple four primitive-cell layer model, oriented to expose Pt (111), to initialise the QMC. This stochastic approach solves the Schrödinger equation. It recently came of age for heterogeneous systems involving solids. During hydrolysis of carbon monoxide, initial O-H bond stretch is rate-limiting. Its dissociation energy is offset by surface Pt-H bond formation. The reactive formate (H-O-C=O) species formed by initial hydrolysis of CO, also interacting with a vicinal Pt. The products are carbon dioxide and hydrogen, with a H-atom dissociated from the formate species and another desorbed from Pt(111). This yields 100% hydrogen gas.
@@ 1 Introduction
Here, platinum catalysts are re-visited. We previously studied them for the water-gas shift reaction [3]. The Pt(111) is the most reactive of cleavages, others are relatively inert. It adsorbs carbon-monoxide, without being poisoned by it. The key hydration step was bench-marked using a sophisticated full configuration-interaction (FCI) wave-function for the embedded active-site reaction intermediate complex. Post Hartree-Fock FCI is time-consuming and scales explosively with system size. Therefore, we now correlate ground-state wave-functions directly, using a generic Jastrow-factor, back-flow and a very careful variational optimisation of the fully converged parameter set. This full parameter set is used for both Transition State (TS) and asymptotic structures involving the same atoms. This strategy has the obvious advantage of ensuring the Slater-Jastrow wave-function will be size consistent. The role of CI is taken over by a generic optimised Jastrow factor. This must be expanded to high order, to converge until coefficients are below a tiny threshold (10-14). It is necessary to use the iso-densic approximation, at least insofar as the compared structures for asymptote and TS comprise the same atoms. Failure to do this can lead to minima that are not size-consistent. The major sources of error, finite-size effect and fixed-node error can be made small. The former is specifically reduced when focussing on surface effects, as studied by Rubenstein et al [37]. This work builds on proof that fixed-node error cancels when resolving energy differences between systems of the same atoms, with few rearranged (i.e. TS and asymptote) [4].
@@ 2 Theoretical Method:
Chemical reactions modify electronic correlation energy, at metal catalyst surfaces. Our simple model truncates fcc metals, with a close-packed (111) planar interface. The model compares physisorbed reactants to a transition-state (TS) geometry, located using a Quantum Monte Carlo (QMC) force constant calculation (below). The Transition-State geometry can be optimized using these force constants, as programmed during our earlier work developing the CASINO software [25]. Two structures are defined, involving the same atoms with some bonds modified (water approaching adsorbed CO, on the one-hand, and the TS on the other, for activation barrier calculation). The TS is compact, however, when the water is 7 Å  from the surface (physisorbed), there is a small residual interaction (generally missing from DFT, but accessible to QMC). The TS involves chemisorption that stabilises its product-like form (as a precursor to actual overlap-driven bond formation). A reference active site is defined with a high-level wave-function to start QMC. It caters for reacting molecules and surface atoms describe the metal. This is embedded using periodic potentials. Green’s function methods are applied to heterogeneity of the electronic potential due to the active site [38]. It is input to begin a Quantum Monte Carlo calculation, as follows:
@@
This is already promising, since Density Functional Theory (DFT) often used in these complex systems has been shown to give results opposed to observation. In one example, the reaction appears to be more difficult on the catalyst surface [22]. In other DFT work, as the pseudo-potential improves or the number of electrons approaches the all-electron system, the barrier, initially far below that measured, continues to decrease [23]. These tendencies for DFT results make such an approach unsatisfactory for the bench-marking of barriers. DFT can be used to explore surrounding potential energy surfaces reliably with ab initio or experimental input [24]. Developing and applying QMC calculations on this scale require access to a supercomputer. Since the catalysts are often heavy metals (in this case platinum, Z=78), we need to use pseudo-potentials. Inclusion of semi-core electrons in the valence is certainly necessary when bonding interactions, in the catalytic process involve the transition metal atoms. It also improves test results for platinum [22]. The copper 3d10 shell is dense in the core-region but platinum has its 5d9 6s1 ground-state shell on average further from the core. Pt is thus less prone ’non-locality’ difficulties in pseudo-potentials [25]. Ni has a 4s2 3d8 ground-state and requires semi-core electrons to reduce the non-locality pitfalls observed for copper[25]. They are limited here by using the Casula algorithm [27]. Comparing asymptote and QMC optimised TS geometries for reaction barriers also involves the same atoms, limiting PP errors. Co-ordination of the TS to the metal surface occurs but the distance between atoms is 40 % more than equilibrium bond–lengths and so the role of semi-core electrons is minor. An argon core for Pt is validated [36].
@@ 2.2.1 Generic Jastrow factor
The generic Jastrow factor used (in Common Algebraic Specification Language (CASL)), had polynomials in e-e, e-n (to order 9) and e-e-n (to order 4), with cut off radius 10 au.
@@
In the present work, DMC runs are short, because the Jastrow factor is fully optimised, and we have obtained convergence for variance followed by energy Variation Monte Carlo (VMC). High-order expansions in the instantaneous inter-particle distance (9 for 2 particle an 4 for 3 particle) lead to large parameter sets (550 here), since individual atoms are treated. When all unbound Pt-atoms are grouped, and also with 3rd order f-expansion, VMC converges faster, but to higher energy. It is therefore an advantage to have converged the complete Jastrow factor, and operate with minimum energies.
@@ 2.3 Diffusion Monte Carlo:
During DMC runs, the ground-state component of the trial wave function is projected out by solving the Schrödinger equation (SWE) in imaginary time. DMC is carried out in the fixed-phase approximation, which uses the nodes from the input trial wave-function. An updated overview of the VMC and DMC methods is given in [19]. These trial wave-functions can be optimised with a complex Jastrow factor, [26] because they potentially provide the input with exact nodes. This improves the single Slater determinant describing a ground-state from the DFT orbitals for heterogeneous systems. DFT ground-state orbitals decay exactly, but Jastrows are needed to improve its nodes. They are also improved by embedding the MRCI active site [10].
@@ 3 Models
One author consistently used Quantum Monte Carlo methods at chemical accuracy ( of 1 kcal/mol = 4.2 kJ/mol) for activation barriers of reactions at metal catalyst surfaces, since publishing test benchmarks on hydrogen dissociation on Cu(111) in 2015. Progress has been made in controlling error sources with the Casino QMC code for periodic systems (Doblhoff-Dier and Hoggan [25, 20]). Our recent publication gives the associated rate-limiting barrier to within 1kJ/mol [10, 9]. This benchmark uses a Multi-reference (MRCI) trial wave-function, suited to bond breaking and formation and an embedded active site approach to the catalyst, allowing high-level input to be used for a small molecular active site, before the periodic QMC. Other applications of QMC to catalysis and weak interactions of accuracy close to 1 kJ/mol are rare and often used to calibrate DFT functionals for cheaper investigations [12, 13, 14]. The above reference asymptotic guiding wavefunction is obtained for the structure below, and simply taken from our previous work [10, 9]. Reliable DFT results on catalytic hydrogen production are [17, 18]. In this work, we compare the TS, from single determinant input to the multi-reference benchmark.
Figure 2: Primitive unit cell for Ni, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  TS closeup (note H-atom at x=y=0.
@@ 3.1 The stages of our approach are as follows:
2-Resulting Slater determinant wave-functions are then expressed in Slater-Jastrow form (Kohn-Sham determinant and Jastrow function product).
3-This trial wave function contains arbitrary parameters (polynomial coefficients of inter-particle distances in the Jastrow factor) that are optimised using Variational Monte Carlo (VMC).
4-The resulting optimised trial wave function is used to generate initial configurations variationally for the Diffusion Monte Carlo step (DMC). Pseudo-potentials have been discussed in previous work [25] and require a brief recap in the section below.
@@ 4 Results
The asymptote is taken from our benchmark, since it is the TS that can be expected to show strong multi-reference character, and we test the infuence of this on activation barrier accuracy by comparing it with single-determinant work conducted in the present study. Here, only $4000$ DMC steps for each geometry converge it, to an error in standard error (se) ratio $R<1/10$. Furthermore, the grid-sizes used are modest (2 2 1), compared with converged DFT grids of 16 16 1. Converged DFT total energies are used as control-variate in the QMC twist-averaging process [19], using 4 twists at present and time-step $\tau$ = 0.005 au.
Defining ”corr” as time-correlation, given as a number of steps. E in au. $4200$ steps; corr=262 use the best reblock length (256) standard error (se) = 1.583 mHa/ error in se 0.3066 mHa
Ets = $-462.035595$ approx. -462.0356 $\bf(max-462.0340)$
$2770$ steps; corr gives $32$ 1.03/32 Ets = -462.03324 $\bf(max-462.03424$   similar to above)
Best
$-462.0356\pm 1.6mHa$
Asymptote (referenvce value) [10]
$-462.0629\pm 0.2mHa$
BARRIER = $71.2$ kJ/mol ±$\pm 4.2$ kJ/mol This model reaches the threshold of chemical accuracy, in spite of limited resources required (1 kcal/mol = 4.18 kJ/mol). The error bars on the TS are 5 times those obtained using MRCI and FCI embedded active site input, i.e. compared to the benchmark [10]. Nevertheless, the activation barrier agrees with the 71 ±$\pm 0.7$ kJ/mol obtained in the benchmark. They are therefore consistent, but could be several kJ/mol apart.
@@
A further run, extrapolated to the zero time-step (tau, $\tau\rightarrow 0$) limit converges well, with a correlation length just over 64 steps (reblocked as 64), for which the accuracy of the barrier is $\pm$ 2kJ/mol. This is continued for a total of 6000 steps and the onset of long-range oscillatory behaviour is observed, leading to an amplitude of 1.5 kJ/mol which is within the variance obtained by twist-averaging over 4 twists. $5600$ steps; corr=82 use best block (64) $64$ converged se is 0.93 mHa (error in standard error 0.0735 mHa), less than R = 1/10 Ets = $-462.036720$ (max -462.0360) absolute difference with $\tau$ = 0.005 runs is 3kJ/mol or 2 standard errors, so not beyond Gaussian significance.
taking the reference asymptote: EXTRAPOLATED $\tau$ BARRIER: = $68.8$ kJ/mol ±$\pm 2.2$ kJ/mol Time-step bias is removed by the extrapolation of steps to zero. Accuracy is improved, by a factor two, remaining more than twice that of the benchmark. All thtree barrier estimates are well within 2kJ/mol and the ’extrapolated tau’ mark has a variationally lower-better TS). The barrier agrees with the 71kJ/mol obtained in the benchmark in [10], however, it has twice the error bar. They are consistent (up to a couple of kJ/mol apart). Finally, we can quote the activation barrier, with a significantly narrowed 95 % confidence interval (which also betters the long calculations of [15]) as: Ea = 70.2 kJ/mol ±$\pm 2.2$ kJ/mol This is consistent with all the numerical results of this work, and close to the benchmark at 71 ±$\pm 0.7$ kJ/mol for the rate-limiting barrier of CO hydrolysis on Pt(111) remains our reference.
@@
The 4 mesh thick slab of section 3, co-ordinates adsorbed molecules above ithe (111) face. It has also become clear that these DMC runs include a sytematic error, also related to the 4-mesh Platinum catalyst model. It stems from the Pt-pseudo-potential that has significant non-locality [25] and was contained by high order non-local grid integration. Convergence tests show non-local PP error in total energy is $\leq 1.5$ kJ/mol.
@@
Despite this, DMC populations of walkers fluctuate by at most 5-10 % and the R threshold for converged runs at 1/10 is reached by 5 600 data-points without significant time-step bias and blocks of 256 points in averaging. Fixed-node error is also present, however it can be made small, even with the present Single-determinant wave-functions [26]. This is represents a metal surface and the reaction is conducted realistically. the bulk solid underneath is periodic.

### v3

@@ Abstract
Many scientists are now researching hydrogen production, because it provides a clean, sustainable alternative to fossil fuel [4]. Research work and prototyping various aspects of hydrogen power continues to advance.
Over 90 % of all chemical manufacture uses a solid catalyst. This work describes catalytic selective hydrogen production (hydration of pre-adsorbed CO, during the water-gas shiftreaction) optimising reactant structure on a model catalyst. Focus is on partial O-H bond dissociation after carbon-monoxide is co-adsorbed with water at Pt(111). Finally, hydrogen gas is given off. Many chemical reactions involve bond-dissociation. This process is often the key to rate-limiting reaction steps at solid surfaces. Bond-breaking is poorly described by Hartree-Fock and DFT methods, our embedded active site approach is used. We showcase Quantum Monte Carlo (QMC) methodology using the ground-state Slater Determinant of a simple four primitive-cell layer model, oriented to expose Pt (111), to initialise the QMC. This stochastic approach solves the Schrödinger equation. It recently came of age for heterogeneous systems involving solids. During hydrolysis of carbon monoxide, initial O-H bond stretch is rate-limiting. Its dissociation energy is offset by surface Pt-H bond formation. The reactive formate (H-O-C=O) species is formed, and interacts with a vicinal Pt. The products are carbon dioxide and hydrogen, with a H-atom dissociated from the formate species and another desorbed from Pt(111). This yields 100% hydrogen gas.
@@ 1 Introduction
Here, platinum catalysts are re-visited. We previously studied them for the water-gas shift reaction [3]. The Pt(111) is the most reactive of cleavages, others are relatively inert. It adsorbs carbon-monoxide, without being poisoned by it. The key hydration step was bench-marked using a sophisticated full configuration-interaction (FCI) wave-function for the embedded active-site reaction intermediate complex. Accurate post Hartree-Fock FCI is time-consuming and scales explosively with system size. Therefore, we now correlate ground-state wave-functions directly, using a generic Jastrow-factor (gJF), back-flow and a very careful variational optimisation of the fully converged parameter set. This full parameter set is used for both Transition State (TS) and asymptotic structures involving the same atoms. The advantage of this strategy lies in ensuring the Slater-Jastrow wave-function will be size consistent. The role of CI is taken over by a generic optimised gJF. The gJF must be expanded to high order, to converge until coefficients smaller than, say (10-14). It is necessary to use the iso-densic approximation, at least insofar as the compared structures for asymptote and TS comprise the same atoms. Failure to do this can lead to minima that are not size-consistent.
Main sources of (non-statistical) error, finite-size effect and fixed-node error can be made small. The former is specifically reduced when focussing on surface effects, as shown by Rubenstein et al [37]. We use the fact that fixed-node error cancels when evaluating energy differences between systems comprising the same atoms, with few rearranged (TS and asymptote) [4].
@@ 2 Theoretical Method:
Chemical reactions modify electronic correlation energy, at metal catalyst surfaces. Fcc metals are truncated with a close-packed (111) planar interface. The model compares physisorbed reactants to a transition-state (TS) geometry, located using a Quantum Monte Carlo (QMC) force constant calculation. Transition-State nuclear positions can be optimized using these force constants, we programmed for the CASINO software [25]. QMC is then applied to the electrons, for ’clamped’ nuclei. Two structures are defined, involving the same atoms with some bonds modified (water approaching adsorbed CO, and the hydration TS, for activation barrier calculation). The TS is compact, however, when the water is 7 Å  from the surface (physisorbed), there is a small residual interaction (generally missing from DFT, but accessible to QMC). The TS involves chemisorption that stabilises its product-like form (as a precursor to actual overlap-driven bond formation). A reference active site is defined with a high-level wave-function to start QMC. It caters for reacting molecules and surface atoms describe the metal. This is embedded using periodic potentials. Green’s function methods are applied to heterogeneity of the electronic potential due to the active site [38]. It is input to begin a Quantum Monte Carlo calculation, as follows:
@@
This is already promising, since Density Functional Theory (DFT) often used in these complex systems has been shown to give results opposed to observation. In one example, the reaction appears to be more difficult on the catalyst surface [22]. In other DFT work, as the pseudo-potential improves or the number of electrons approaches the all-electron system, the barrier, initially far below that measured, continues to decrease [23]. These tendencies for DFT results make such an approach unsatisfactory for the bench-marking of barriers. DFT can be used to explore surrounding potential energy surfaces reliably with ab initio or experimental input [24]. Developing and applying QMC calculations on this scale require access to a supercomputer. Since the catalysts are often heavy metals (in this case platinum, Z=78), we need to use pseudo-potentials. Inclusion of semi-core electrons in the valence is certainly necessary when bonding interactions, in the catalytic process involve the transition metal atoms. It also improves test results for platinum [22]. The copper 3d10 shell is dense in the core-region and delocalised in the conduction bands, but platinum has its 5d9 6s1 ground-state shell on average further from the core. Cu and Ni (with a 4s2 3d8 ground-state) are prone to ’non-locality’ effects, in pseudo-potentials [25]. Pt is less affected. All require semi-core valence shells [25]. These effects are limited by the Casula T-move algorithm [27] (in DMC). Comparing asymptote and QMC optimised TS geometries for reaction barriers also involves the same atoms, limiting PP errors. Co-ordination of the TS to the metal surface occurs but the distance between atoms is 40 % more than equilibrium bond–lengths and so the role of semi-core electrons is minor. An argon core for Pt is validated [36].
@@ 2.2.1 Generic Jastrow factor
The generic Jastrow factor used (in Common Algebraic Specification Language (CASL)), comprised polynomials in e-e, e-n (to order 9) and e-e-n (to order 4), with cut off radius around 10 a.u. (see aapendix).
@@
In the present work, DMC runs require few data-points, the Jastrow factor being fully optimised, and we have obtained convergence for variance followed by energy Variation Monte Carlo (VMC). High-order expansions in the instantaneous inter-particle distance (9 for 2 particle an 4 for 3 particle) give large parameter sets (550 here with individual atoms are treated). When all unbound Pt-atoms are grouped, and also with 3rd order f-expansion, VMC converges faster, but to higher energy. It is therefore an advantage to have converged the complete Jastrow factor, and operate with minimum energies.
@@ 2.3 Diffusion Monte Carlo:
During DMC runs, the ground-state component of the trial wave function is projected out by solving the Schrödinger equation (SWE) in imaginary time. DMC is carried out in the fixed-phase approximation, which uses the nodes from the input trial wave-function. An updated overview of the VMC and DMC methods is given in [19]. These trial wave-functions can be optimised with a complex Jastrow factor, [26] because they potentially provide the input with exact nodes. This improves the single Slater determinant describing a ground-state from the DFT orbitals for heterogeneous systems. DFT ground-state orbitals decay exactly, but Jastrows are needed to improve its nodes. They are also improved by embedding the Multi-reference (MRCI) active site [10].
@@ 3 Models
One author consistently used Quantum Monte Carlo methods at chemical accuracy ( of 1 kcal/mol = 4.2 kJ/mol) for activation barriers of reactions at metal catalyst surfaces, since publishing test benchmarks on hydrogen dissociation on Cu(111) in 2015. Progress has been made in controlling error sources with the Casino QMC code for periodic systems (Doblhoff-Dier and Hoggan [25, 20]). Our recent publication gives the associated rate-limiting barrier to within 1kJ/mol [10, 9]. This benchmark uses a Multi-reference (MRCI) trial wave-function, suited to bond breaking and formation and an embedded active site approach to the catalyst, allowing high-level input to be used for a small molecular active site, before the periodic QMC. Other applications of QMC to catalysis and weak interactions of accuracy close to 1 kJ/mol are rare and often used to calibrate DFT functionals for cheaper investigations [12, 13, 14]. The above reference asymptotic guiding wavefunction is obtained for the structure in Figure 1, and simply taken from our previous work [10, 9]. Reliable DFT results on catalytic hydrogen production are [17, 18]. In this work, we compare the TS (the strcture is optimised with QMC forces and plotted in Figure 2), from single determinant input to the multi-reference benchmark.
Figure 2: Primitive unit cell for Ni, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  TS closeup (note H-atom at x=y=0.)
@@ 3.1 The stages of our approach are as follows:
3-Resulting Slater determinant wave-functions are then expressed in Slater-Jastrow form (Kohn-Sham determinant and Jastrow function product).
4-This trial wave function contains arbitrary parameters (polynomial coefficients of inter-particle distances in the Jastrow factor) that are optimised using Variational Monte Carlo (VMC).
5-The resulting optimised trial wave function is used to generate initial configurations variationally for the Diffusion Monte Carlo step (DMC). Pseudo-potentials have been discussed in previous work [25] and require a brief recap in the section below.
@@ 4 Results
To test the validity of activation barriers from QMC with single-determinant input, the asymptote is taken from our benchmark, since it is the TS that can be expected to show strong multi-reference character. The benchmark activation barriers are compared with single-determinant work conducted in the present study. Here, only $4000$ DMC steps for each geometry converge it, to an error in standard error (se) ratio $R<1/10$. Furthermore, the grid-sizes used are modest (2 2 1), compared with converged DFT grids of 16 16 1. Converged DFT energy values are the control-variate in our QMC twist-averaging process [19], using 4 twists here, and time-step $\tau$ = 0.005 au. Tabulated averaged raw QMC runs (a.u.): Defining ”corr” (time-correlation) as a number of steps. The best reblock length, is labeled ’rl’. The standard error (se) and E, are in Hartree (Ha). Steps corr best rl se mHa error in se Ets a $4200$ 262 256 1.583 0.3066 $-462.035595$ b $2770$ 34 32 1.03 0.32 -462.03324 Tabulated analysis of error-strewn barrier height, in Hartree (Ha) and standard error (se) units specified: Run. Etot maximum se mHa Barrier kJ/mol se kJ/mol a $-462.0356$ $-462.0340$ 2 - - b $-462.0332$ $-462.03424$ 2 - - best $-462.0356$ 1.6 - - asymptote $-462.0629$ 0.2 $71.2$ $4.2$ benchmark - - 0.2 $71$ $0.7$
@@
This model reaches the threshold of chemical accuracy, in spite of limited resources required (1 kcal/mol = 4.18 kJ/mol). The TS standard error is 5 times that of the benchmark [10], obtained using MRCI (FCI) embedded active site input. Nevertheless, the activation barrier agrees with the 71 ±$\pm 0.7$ kJ/mol obtained in the benchmark. They are therefore consistent, but could be several kJ/mol apart. A further run, extrapolated to the zero time-step (tau, $\tau\rightarrow 0$) limit converges well, with a correlation length just over 64 steps (reblocked as 64), for which the accuracy of the barrier is $\pm$ 2kJ/mol. This is continued for a total of 6000 steps and the onset of long-range oscillatory behaviour is observed, leading to an amplitude of 1.5 kJ/mol which is within the variance obtained by twist-averaging over 4 twists. Time-step bias is removed by the extrapolation of steps to zero. Accuracy is improved, by a factor two, remaining more than twice that of the benchmark. All thtree barrier estimates are well within 2kJ/mol and the ’extrapolated tau’ mark has a variationally lower-better TS). The barrier agrees with the 71kJ/mol obtained in the benchmark in [10], however, it has twice the error bar. They are consistent (up to a couple of kJ/mol apart).
After longer runs and extrapolation to zero time-step, we again take the reference asymptote. The absolute difference with $\tau$ = 0.005 runs is 3kJ/mol or 2 standard errors, so not beyond Gaussian significance. Tabulated barriers from longer runs (refined barrier heights) Again, ”corr” is time-correlation, given as a number of steps. The best reblock length is labeled ’rl’. The se (standard error) and E, are in Hartree steps Ets max. corr/rl se mHa err. in se Ea kJ/mol a $5600$ $-462.03672$462.036 84 / 64 0.93 0.0735 $68.8$$\pm 2.2$
This is consistent with all the numerical results of this work, and close to the benchmark at 71 ±$\pm 0.7$ kJ/mol for the rate-limiting barrier of CO hydrolysis on Pt(111) remains our reference.
@@
Finally, we can quote the activation barrier, with a significantly narrowed 95 % confidence interval (which also betters the long calculations of [15]) as: Ea = 70.2 kJ/mol ±$\pm 2.2$ kJ/mol The 4 mesh thick slab (see section 3) has adsorbed molecules above the (111) face. These DMC runs include sytematic error, due to the 4-mesh Platinum catalyst model. It stems significant non-locality [25] in the Pt pseudopotentil, and was contained by high order non-local grid integration. Convergence tests show non-local PP error in total energy is $\leq 1.5$ kJ/mol.
@@
Having shown this ’non-locality’ PP-defect persists in the DMC runs (gently oscillating energy values), causing the populations of its walkers fluctuate by at most 5-10 % and the R threshold for converged runs at 1/10 is reached by 5 600 data-points without significant time-step bias and blocks of 256 points in averaging. Fixed-node error is also present, however it can be made small, even with the present Single-determinant wave-functions [26]. This is represents a metal surface and the reaction is conducted realistically. the bulk solid underneath is periodic.